A certain sum amounts to Rs. 4800 in 4years and to5250 in 5 years at simple interest . Find the interest in 2 years? Please sir help me....

Grade:Select Grade

1 Answers

Vikas TU
6 years ago
Dear Student,
 let principal amount be p
       4800=P+PR(4)/100
       5250=P+PR(5)/100
Solve these two equations to get PR= 45000
Interest in 2 years = PRT/100 = 45000*2/100= 900.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)